carmelle:

Factor completely: 32x^2 − 80x + 50

carmelle:

A. 2(4x − 5)(4x − 5) B. 2(4x + 5)(4x + 5) C. 2(8x − 5)(8x − 5) D. 2(8x + 5)(8x + 5)

snowflake0531:

Can you divide the whole equation by 2 first?

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 Can you divide the whole equation by 2 first? \(\color{#0cbb34}{\text{End of Quote}}\) yes?

snowflake0531:

So what would you get

snowflake0531:

\[\frac{ 32x^2-80x+50 }{ 2 }\]

carmelle:

im confused :(

snowflake0531:

We can expand on ^^ equation to \[\frac{ 32x^2 }{ 2 } - \frac{ 80x }{ 2 } + \frac{ 50 }{ 2 }\]

carmelle:

ok, 16x, 40x and 25?

snowflake0531:

Yes so we have \[16x^2 -40x + 25\]

snowflake0531:

Now, which ones can you eliminate looking at 16x^2

snowflake0531:

hint, look at the choices where the differences are 4 and 8

bhaddienene08:

common factor 32x-80+50

carmelle:

we can eliminate a and b?

snowflake0531:

No, if we eliminated a and b, we would be left with c and d, which would maek 64x^2

bhaddienene08:

2(16X-40+25

carmelle:

ohhhh

bhaddienene08:

Use the sum-product pattern Common factor from the two pairs

snowflake0531:

So you eliminate the third and fourth

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 So you eliminate the third and fourth \(\color{#0cbb34}{\text{End of Quote}}\) ok

snowflake0531:

And so now, we just look at the middle one In this case, it is a negative number So, which one would we choose

bhaddienene08:

2(4𝑥−5)^2

carmelle:

It's A

bhaddienene08:

i think it's A

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle It's A \(\color{#0cbb34}{\text{End of Quote}}\) yes

carmelle:

tysm, im so bad at factoring o-o

snowflake0531:

tbh, I usually just like to guess around xd, and not actually think thru it

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 tbh, I usually just like to guess around xd, and not actually think thru it \(\color{#0cbb34}{\text{End of Quote}}\) lol, I'll try that

snowflake0531:

So, especially with these, you have choices, you could just factor it out using FOIL and see which one is right

carmelle:

ok :D